BRICS Notes Series, Abstracts, 2000

July 7, 2003

This document is also available as PostScript and DVI.

References

[NS-00-8] PostScript, PDF.

Anders Mller and Michael I. Schwartzbach. The XML Revolution. December 2000. 149 pp. Abstract: This 130+ page slide collection provides an introduc-

tion and overview of XML, Namespaces, XLink, XPointer, XPath, DSD, XSLT, and XML-QL, including selected links to more information about each topic.

[NS-00-7] PostScript, PDF, DVI.

Nils Klarlund, Anders Mller, and Michael I. Schwartzbach. Document Structure Description 1.0. December 2000. 40 pp. Abstract: Document Structure Description 1.0 is a complete spec-

ification of a new XML notation for describing classes of XML documents. The notation is designed to be a simple tool based on familiar concepts. DSDs provide more flexible and precise structural descriptions than possible with DTDs or the current XML Schema proposal. A DSD generates a CSS-like default mechanism independent of formatting models. Finally, it allows an extension mechanism so that DSDs may be updated with new structural concepts.

[NS-00-6] PostScript, PDF.

Peter D. Mosses and Hermano Perrelli de Moura, editors. Proceedings of the Third International Workshop on Action Seman-

tics, AS 2000, (Recife, Brazil, May 15–16, 2000), August 2000. viii+148 pp. Abstract: Action Semantics is a practical framework for formal se-

mantic description of programming languages. Since its appearance in 1992, Action Semantics has been used to describe major languages such as Pascal, SML, ANDF, and Java, and various tools for processing action-semantic descriptions have been developed. The AS 2000 workshop included reports of recent achievements with the foundations and applications of Action Semantics, presentations and demonstrations of tool support for actionsemantic descriptions, and discussion of a proposal for a new (and significantly simpler) version of Action Notation.

[NS-00-5] PostScript, PDF. Claus Brabrand. <bigwig> Version 1.3 — Tutorial. September 2000. ii+92 pp. Abstract: This tutorial to the <bigwig>-language is based on a num-

ber of example programs illustrating the features of the language.

Contents • Introduction

• Getting Started

• Dynamic Documents

• PowerForms

• Database

• Concurrency Control

• Macros .

[NS-00-4] PostScript, PDF. Claus Brabrand. <bigwig> Version 1.3 — Reference Manual. September 2000. ii+56 pp. Abstract: This reference manual concisely describes the whole

<bigwig>-language.

Contents • Keyword Index

• Lexical Structure

• Syntax

• Operators

• Scope Rules

• Types

• Core Language

• Control Structures

• Files

• Macros

• Form Input

• Formats • Dynamic Documents

• Database

• Security

• Concurrency Control

• Web Specifics

• Time

.

[NS-00-3] PostScript, PDF.

Patrick Cousot, Eric Goubault, Jeremy Gunawardena, Maurice Her-

lihy, Martin Rauen, and Vladimiro Sassone, editors. Preliminary Proceedings of the Workshop on Geometry and Topol-

ogy in Concurrency Theory, GETCO ’00, (State College, USA, August 21, 2000), August 2000.

vi+116 pp. Abstract: This volume contains the preliminary proceedings of

GETCO ’00, held in State College (Pennsylvania, USA) on August 21, 2000. The workshop aims at bringing together researchers from both the mathematical (geometry, topology, algebraic topology etc.) and computer scientific side (concurrency theorists, semanticians, researchers in distributed systems etc.) with an active interest in these or related developments. The final proceedings will appear as volume 39 in the ENTCS series, which can be found at www.elsevier.nl/locate/entcs/volume39.html.

[NS-00-2] PostScript, PDF. Luca Aceto and Bjrn Victor, editors. Preliminary Proceedings of the 7th International Workshop on Ex-

pressiveness in Concurrency, EXPRESS ’00, (State College, Pennsylvania, USA, August 21, 2000), August 2000.

vi+130 pp. Abstract: This volume contains the preliminary proceedings of

EXPRESS’00, which was held in State College (Pennsylvania, USA) on 21 August 2000. It includes the six papers that were selected for presentation by the program committee, together with the contribution by the invited speaker, Neil D. Jones (DIKU, Denmark). The final proceedings will appear as volume 39 in the ENTCS series, which can be found at www.elsevier.nl/locate/entcs/volume39.html.

[NS-00-1] PostScript, PDF. Bernd Grtner. Randomization and Abstraction — Useful Tools for Optimization. February 2000. 106 pp. Abstract: This report contains notes for a course on randomized

methods in geometric optimization, held first at BRICS, Aarhus University, Denmark and later at the Institute for Computer Science, ETH Zrich, Switzerland.